Question 3: National Accounting (20 Marks) In Table 2 you find figures taken from some Countrys national accounts for the 2018 fiscal year. Figures are in billions of dollars. Find: a) Savings b) Private Savings c) Net Government Income d) Gross National Product e) Net Exports f) Government Interest Payments g) Current Account Balance
Variable Y (GDP) NFP (Net Factor Payments) C (Consumption) G (Government Expenditures) PDI (Personal Disposable Income) I (Investment) T (Taxes) TR (Government Transfers) Value 2000 100 1400 200 1600 160 640 130
Answer #1

a) Saving = Y-C-G = 2000-1400-200 = 400
b) Private Saving = (Y-T)-C = (2000-640)-1400 = -40
c) Net government Income = T-TR-INT = 640-130-10=500
d) Gross national product = Y+NFP = 2000+100=2100
e) Net exports = Y-C-I-G =2000-1400-160-200 =240
f) Government interest payments
PDI = Y+NFP+TR+INT-T
INT = PDI-Y-NFP-TR+T = 1600-2000-100-130+640 = 10
g) Current account balance = NFP+NX = 100+240 = 340